OpenStudy (anonymous):

Harry can rake the leaves in the yard 8 hours faster than his little brother Jimmy can. If they work together, they can complete the job in 3 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jimmy to complete this job on his own.

OpenStudy (anonymous):

triciaal

OpenStudy (triciaal):

@Pico33 please help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok so Harry can rake the leaves in the yard 8 hours faster than his little brother Jimmy can. If they work together, they can complete the job in 3 hours.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x(x-8)/(x+x-8) = 3

OpenStudy (anonymous):

Could you walk me through it? im not sure how to work through it.

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

thanks :)

OpenStudy (triciaal):

@mathstudent55?

OpenStudy (anonymous):

X is Jimmy's time if he completes the job alone. x(x-8)/(x+x-8) = 3 (x^2 - 8x)/(2x-8) = 3 cross multiply then simplify and solve for x

OpenStudy (anonymous):

so what do you get?

OpenStudy (anonymous):

could we add together the 8x and the 2x?

OpenStudy (anonymous):

x^2-8x / 2x - 8 = ?

OpenStudy (anonymous):

3/1

OpenStudy (anonymous):

im not sure what to do next

OpenStudy (anonymous):

x^2-8x/6x-24

OpenStudy (anonymous):

x^2-14x+24=0

OpenStudy (anonymous):

factor it (x-12)(x-2)=0

OpenStudy (anonymous):

x = 12

OpenStudy (anonymous):

Jimmy = 12 hours Harry = 4 hours

OpenStudy (anonymous):

So do you see that Harry is 8 hours faster than his brother?

OpenStudy (anonymous):

so your answer is 12

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

@icefire @triciaal

OpenStudy (anonymous):

yes thank you

OpenStudy (triciaal):

@icefire do you understand how we got the equation to start? @Pico33 thank you so much because I am exhausted.

OpenStudy (triciaal):

it's almost 3am I should go to sleep

OpenStudy (triciaal):

thank you